# Scalp Diagnostic System With Label-Free Segmentation and Training-Free Image Translation

### v2

@@ Abstract
Scalp diseases and alopecia affect millions of people around the world, underscoring the urgent need for early diagnosis and management of the disease. However, the development of a comprehensive AI-based diagnosis system encompassing these conditions remains an underexplored domain due to the challenges associated with data imbalance and the costly nature of labeling. To address these issues, we propose “ScalpVision”, an AI-driven system for the holistic diagnosis of scalp diseases and alopecia.
In ScalpVision, effective hair segmentation is achieved using pseudo image-label pairs and an innovative prompting method in the absence of traditional hair masking labels. This approach is crucial for extracting key features such as hair thickness and count, which are then used to assess alopecia severity. Additionally, ScalpVision introduces DiffuseIT-M, a generative model adept at dataset augmentation while maintaining hair information, facilitating improved predictions of scalp disease severity. Our experimental results affirm ScalpVision’s efficiency in diagnosing a variety of scalp conditions and alopecia, showcasing its potential as a valuable tool in dermatological care.
@@ 1 Introduction
Approximately 80 million people in the United States are affected by male and female pattern hair loss [36]. Additionally, a study by Elewski et al. [14] reveals that nearly $90\%$ of adults in the U.S. experience various scalp disorders. Among these, alopecia is especially common, impacting a significant portion of the population. The causes of alopecia are diverse, including aging, genetics, stress, and overall health conditions [46]. Therefore, early diagnosis is key to preventing the progression of these scalp-related diseases [35, 37].
Recognizing the importance of early detection, numerous studies have explored scalp disease diagnosis using microscopic scalp imagery. For example, Chang et al. [7] proposed a scalp diagnostic system, Kim et al. [24] concentrated on follicle detection, and Seo et al. [43] identified features of hair loss for diagnosing alopecia. However, these studies often treat scalp and hair issues separately, overlooking a more integrated diagnostic approach.
Figure 1: Data distribution and examples of AI-Hub microscopic scalp imaging training dataset. The images on the left side represent examples of each scalp disease condition at different severity and the charts on the right side represent data distribution of different severity within each scalp condition.
@@
Figure 2: ScalpVision pipeline overview. Here, $I$ denotes the original image. The hair segmentation mask generated by the segmentation model $S$, which is trained using a pseudo-training set, is represented as $\hat{M}$. The mask produced by the SAM is denoted as $M_{\text{AP}}$, and $M$ signifies the final, combined hair segmentation mask. The “Automatic Prompt”, used for refining segmentation, is derived from $\hat{M}$.
Acquiring comprehensive labels for medical diagnostic systems presents a significant challenge in medical research. This difficulty primarily arises from the need for specialized domain knowledge in image annotation, which significantly increases labeling costs. Furthermore, data imbalance is a frequent issue, largely due to biases in the training process. For instance, most datasets focusing on scalp conditions disproportionately represent healthy scalps, leading to a scarcity of data on severe scalp diseases and alopecia. As a result, there are no publicly accessible datasets that offer detailed labels for both scalp conditions and alopecia in scalp images.
Despite the lack of publicly available datasets, we accessed a specialized dataset from AI-hub [1] for classifying the severity of scalp conditions and alopecia  [^1] This dataset is provided by ‘The Open AI Dataset Project (AI-Hub, S. Korea)’ and is exempt from IRB approval as it does not contain any information that can identify individuals.. This dataset categorizes the severity of scalp conditions, including dandruff, excess sebum, and erythema, into four levels (good, mild, moderate, severe) across 96K images. It also includes about 24K images for alopecia severity classification. The distribution of this dataset’s attributes is illustrated in Figure 1. However, this dataset lacks detailed information on hair and follicles, which is crucial for more in-depth diagnostic analyses.
To address these challenges, we developed “ScalpVision,” a comprehensive system designed for in-depth assessment of scalp health, as illustrated in Figure 2. ScalpVision’s pipeline integrates three key components: 1) hair segmentation, 2) diagnosis of scalp conditions, and 3) prediction of alopecia severity. The backbone of ScalpVision is its advanced segmentation technique. This technique employs label-free learning and zero-shot segmentation to effectively discern and extract hair-related features. It combines masks generated by a model trained on synthetic image-label pairs with those from the point-guided Segment Anything Model (SAM) [26]. For diagnosing scalp conditions, we leverage DiffuseIT-M, a diffusion-based image-to-image translation model, adept at addressing data imbalance. Finally, in the alopecia severity assessment phase, we utilize hair masks to derive features like hair thickness and count, which are then analyzed using machine learning models to provide accurate evaluations.
In summary, our main contributions are as follows:
  • We present a comprehensive diagnostic pipeline for microscopic scalp imagery that assesses both scalp condition and alopecia severity.
  • We develop a novel segmentation approach, combining heuristic-driven pseudo-labels with SAM to effectively segment images in the absence of ground truth labels.
  • We introduce DiffuseIT-M, a mask-guided, diffusion-based image-to-image translation method, facilitating targeted control over specific image regions without additional training.
We contribute to the field of scalp diagnosis by sharing our code on GitHub  [^2] https://github.com/winston1214/ScalpVision.
2 Related Works
Scalp image analysis. In the field of scalp image analysis, hair is a pivotal feature, often utilized for segmentation purposes [44, 48]. Investigations into the number [52, 41] and thickness [23] of hairs on microscopic scalp images have been conducted. Moreover, various approaches utilize microscopic scalp images to diagnose diseases. Systems for classifying scalp diseases and their severity have been developed [7, 25], while others assess alopecia severity using scalp information [24, 43].
Image augmentation. Previous studies have proposed methods for data augmentation by merging images [53, 54] or losing partial information of images [58, 8]. Recent developments in generative models, particularly the diffusion models [20, 40], have led to the proposal of image augmentation methods using synthetic images. Azizi and Trabucco et al. [3, 47] show diffusion models can generate photo-realistic images with given conditioning variables and improve image classification accuracy, even in medical applications [11, 57].
Label-free image segmentation on medical image. The challenges and expenses associated with acquiring medical image data have spurred the development of data-efficient training methods [9] and label-free segmentation models. Kim et al. [22] introduced a synthetic angiogram based on fractal design as a label, training a diffusion module for vessel segmentation. Large pre-trained segmentation models like the Segment Anything Model (SAM) [26] have garnered interest and demonstrated promise across various vision tasks [59]. While these models are increasingly applied to specific medical imaging challenges, success has been limited [17].
3 Our Approach: ScalpVision
@@
To address the challenges of limited and imbalanced datasets in scalp imaging, we introduce ScalpVision, a diagnostic system designed for accurately assessing the severity of scalp-related diseases and alopecia. Central to ScalpVision is a hair segmentation module (Section 3.1), pivotal for both alopecia severity prediction (Section 3.2) and generating diverse scalp images to augment training datasets for scalp condition classification (Section 3.3). The overall pipeline of ScalpVision is illustrated in Figure 2.
@@ 3.1 Label-Free Hair Segmentation
For the precise diagnosis of alopecia and scalp conditions, our initial step involves segmenting hair within microscopic scalp images. However, as the dataset does not provide explicit segmentation labels, conventional supervised learning methods are not feasible.
Heuristic-driven pseudo-labeling. To address this, we first generate pseudo labels for training our segmentation model (S as shown in Figure 2) using prior knowledge. With the intuition that the hair on the microscopic scalp images follows either a linear function or a power function, we generate synthetic images to effectively guide the model to learn hair patterns on the scalp images. For each disease condition, we randomly select one image representing each distinct severity level, extract three smaller patches from regions of the scalp with no visible hair, and draw curves to simulate hair patterns. Additionally, to simulate dandruff noise, circular white shapes are added to these patches, but are not indicated in the pseudo masks, thus training the model to interpret them as noise. We generate $3{,}000$ pseudo-images and corresponding pseudo mask labels, using them to train the U2-Net [39]. This training generates the binary mask, represented as:
$\hat{M}=[\hat{M}(i,j)]\in\{0,1\}^{H\times W},$ (1)
where $H$ and $W$ are the height and width of the image, and $i\in[1,H]$, $j\in[1,W]$ denote pixel coordinates.
@@
Input: Mask $\hat{M}$, bounding box size $n$, and cross-shaped structuring element kernel
Output: Representative hair points from mask $\hat{\mathcal{C}}$
1:  $\mathcal{H}_{\text{copy}}$leftarrow $\hat{{M}}$
2:  ^$\hat{\mathcal{H}}_{\text{skel}}$leftarrow zero array with same size as $\mathcal{H}_{\text{copy}}$
3:  while $\mathcal{H}_{\text{copy}}$$\neq$00 do
4:     Eroded, Dilated $\leftarrow$ MORPHOLOGY(((($\mathcal{H}_{\text{copy}}$, kernel))))
5:     $\hat{K}$leftarrow $\mathcal{H}_{\text{copy}}$- Dilated
6:     ^$\hat{\mathcal{H}}_{\text{skel}}$leftarrow ^$\hat{\mathcal{H}}_{\text{skel}}$lor $\hat{K}$
7:     $\mathcal{H}_{\text{copy}}$leftarrow Eroded
8:  end while
[…]
10:  for all  $(x,y)$in $\mathcal{H}_{\text{skel}}$ do
11:     $\hat{B}$leftarrow $\hat{B}$cup $\{(x-\frac{1}{2}n,y-\frac{1}{2}n,x+\frac{1}{2}n,y+\frac{1}{2}n)\}$
12:  end for
13:  $\hat{B}$leftarrow NMS($\hat{B}$)
[…]
15:  for all $(x_{1},y_{1},x_{2},y_{2})$in $\hat{B}$ do
[…]
17:  end for
18:  return $\hat{\mathcal{C}}$
Automatic prompting for SAM. To refine the hair segmentation mask $\hat{M}$, we utilize the foundation segmentation model, SAM [26], employing a point-prompting method to differentiate hair from scalp without additional training. However, we observed that selecting random points from $\hat{M}$ for positive point prompts often resulted in suboptimal masks. This issue was primarily due to points near the edges of $\hat{M}$ causing confusion for the SAM. Moreover, due to inherent randomness, there were instances in which sampled points clustered in a small localized region, resulting in the SAM segmenting only a limited number of hairs. To address these issues, we have developed an automatic prompting method. This method ensures uniform sampling across $\hat{M}$ and provides high-confidence guidance to the SAM based on the coarse segmentation mask $\hat{M}$. The specific steps of this method are detailed in Algorithm 1.
To extract the distinct features of the hair, we compute the skeletonized mask ^$\hat{\mathcal{H}}_{\text{skel}}$ ($\in\{0,1\}^{H\times W}$) using morphological erosion and dilation following [56]. This method operates by iteratively processing the image, eroding and dilating pixels at the object’s edges until no further pixel removal is possible. Then, we generate bounding boxes around each pixel in ^$\hat{\mathcal{H}}_{\text{skel}}$ with size $n\times n$ where we set $n=10$. These boxes undergo non-maximum suppression (NMS) to filter out the bounding boxes, denoted as $\hat{B}=\{\hat{b}_{j}\}_{j=1}^{k}$, where each box is defined by coordinates $(x_{\text{min}},y_{\text{min}},x_{\text{max}},y_{\text{max}})$. Following this, the mean points of the hair pixels, $\hat{\mathcal{C}}=\{\hat{c}_{j}\}_{j=1}^{k}$, in each bounding box $\hat{B}$ can be determined. For each $\hat{b}_{j}=(x_{1},y_{1},x_{2},y_{2})$, the mean point $\hat{c}_{j}=(\bar{x},\bar{y})$ is given by:
@@
Subsequently, we select positive point prompts for SAM from the calculated mean points $\hat{\mathcal{C}}$. For the negative point prompts, we utilize the inverse of the initial mask, specifically $1-\hat{M}$. These prompts, automatically generated, guide SAM in generating the binary segmentation mask $M_{\text{AP}}$ ($\in\{0,1\}^{H\times W}$).
Mask ensemble. $M_{\text{AP}}$ and $\hat{M}$ complement each other with strengths and weaknesses. $\hat{M}$ is robust against noise like dandruff as it was trained using simulated noise. Meanwhile, $M_{\text{AP}}$, benefiting from SAM’s superior edge detection, excels in constructing a clear boundary between hair and scalp. Therefore, to make a robust hair mask, the final binary mask, $M$, is derived from $\hat{M}$ and $M_{\text{AP}}$ with the AND operation, as follows:
$M=\hat{M}\land M_{\text{AP}}.$ (3)
3.2 Alopecia Severity Prediction
Alopecia primarily manifests through two indicators: hair thinning and follicle reduction [38]. Consequently, hair thickness and count are crucial metrics for assessing its severity. However, lacking ground-truth labels for these features, we have developed straightforward yet effective methods to approximate hair thickness (Figure 3(a)) and hair count (Figure 3(b)). These methods utilize the hair mask $M$ generated earlier, from which we extract bounding boxes $B$, and their respective mean points $C$, following the approach described in Section 3.1.
Figure 3: Feature extraction for alopecia severity prediction.
Hair thickness. Our objective is to measure the thickness of hair within each bounding box in the set $B=\{b_{j}\}_{j=1}^{k}$. Firstly, we define the set of pixel coordinates for hair within a bounding box $b_{j}$ as:
$\mathcal{P}_{j}=\{(x,y)\in b_{j}\mid M(x,y)=1\},$ (4)
where $M(x,y)=1$ indicates the presence of hair at the pixel coordinates $(x,y)$.
Next, we calculate $\mathcal{S}=\{s_{j}\}_{j=1}^{k}$ where each $s_{j}$ represents the slope perpendicular to the hair in the $j$-th bounding box. This slope is derived from the second principal component in a Principal Component Analysis (PCA) of the pixel coordinates $\mathcal{P}_{j}$. A perpendicular line ($\mathcal{L}_{j}$) is then defined using this slope $s_{j}$ and the mean point $c_{j}$:
$\mathcal{L}_{j}=\{(x,y)\mid s_{j}(x-c_{j,x})+c_{j,y}=y\}.$ (5)
Following this, we identify two boundary points of the hair, $p_{j,1}$ and $p_{j,2}$, located where $\mathcal{L}_{j}$ intersects with the edge points of $\mathcal{P}_{j}$. Denoting these edge points as $E(\mathcal{P}_{j})$, the boundary points are formulated as:
$p_{j,1},p_{j,2}=E(\mathcal{P}_{j})\cap\mathcal{L}_{j}.$ (6)
Finally, the pixel-level hair thickness ($t_{j}$) is calculated as the Euclidean distance between these two boundary points:
$t_{j}=||p_{j,1}-p_{j,2}||_{2}.$ (7)
Thus, we obtain a set of hair thicknesses: $\mathcal{T}=\{t_{j}\}_{j=1}^{k}$.
Hair count. To quantify the number of hairs, we utilize the Hough transformation [13], an established technique for line detection through parameter space representation, applied to the hair skeleton $\mathcal{H}_{\text{skel}}$. Inspired by the approach in [45], we enhance the method’s effectiveness for curved hair by implementing the Hough transformation at three different resolutions: original, half, and double. At each of these resolutions, the transformation identifies a set of line segments, where each segment is defined by a pair of endpoints, denoted as $\{(x_{1},y_{1}),(x_{2},y_{2})\}$. The line segments obtained from each resolution level are then resized to correspond to the original resolution scale and collectively merged. During this merging phase, we identify and remove duplicate line segments. This elimination is based on their similarity in vector angle, length, and spatial positioning.
The two features we have predicted, hair thickness and count, not only hold potential for direct use by clinicians in diagnosing alopecia but also present an opportunity for further exploration. Considering the dataset includes labels on the severity of alopecia, we investigate the application of these features as inputs for traditional machine learning models. This approach is aimed at enhancing the prediction of alopecia severity using the available dataset.
Figure 4: Overview of DiffuseIT-M. $x_{src}$ and $x_{trg}$ represent the source image and target image, respectively, with $M$ denoting the mask image of $x_{src}$. In the reverse process, we guide the generation to adhere to the conditions using the objective function. $\odot$ signifies the weighted sum of two images. “DINO” means DINO-ViT.
@@ 3.3 Scalp Condition Classification
Accurately classifying the severity of scalp-related diseases in microscopic images is challenging due to the rarity of extreme cases, as illustrated in Figure 1. To address this issue, we have developed a method that translates given images into scalp images of a desired condition. This method, named DiffuseIT-M, is a diffusion-based image translation model with mask guidance. It is designed to preserve the content (e.g., hair) of the source image while enabling the transformation into various scalp conditions. DiffuseIT-M is based on DiffuseIT [27] and incorporates an image editing technique inspired by Blended diffusion [2] to ensure hair preservation. The overview of DiffuseIT-M is depicted in Figure 4.
Preliminary. In the forward diffusion process, we adopted the DDPM [20] forward process. Given a clean image, $x_{0}\sim q(x_{0})$, Gaussian noise is gradually added at every time steps, $t$, using the Markov chain.
$q(x_{1:T}|x_{0}):=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$ (8)
following the formulations and notations in [33]. This equation can be simplified in terms of $x_{t}$ sampled in one step as follows:
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (9)
where $\epsilon\sim\mathcal{N}(0,\bf{I})$, $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$.
In the reverse diffusion process, $q(x_{t-1}|x_{t})$ is approximated as $p_{\theta}(x_{t-1}|x_{t})$ with the trainable parameter $\theta$. Therefore, it is modeled by a Markov chain with a learned mean and fixed variance, initiated from $p(x_{T})$:
$p_{\theta}(x_{0:T}):=p_{\theta}(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t})$ (10)
where $p_{\theta}(x_{t-1}|x_{t}):=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. Rather than inferring $\mu_{\theta}(x_{t},t)$ directly, it predicts the noise $\epsilon_{\theta}(x_{t},t)$ added to $x_{0}$ to obtain $x_{t}$. Then, ${\mu}_{\theta}(x_{t},t)$ is derived using Bayes’ theorem:
${\mu}_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\Big{)}.$ (11)
Then, the sampling from the reverse diffusion is defined as:
$x_{t-1}={\mu}_{\theta}(x_{t},t)+\sigma_{t}\epsilon.$ (12)
Image translation with mask guidance. To facilitate the transfer of scalp disease characteristics while preserving hair features in our model, we utilize a comprehensive loss function, $\ell_{total}$. This function guides the reverse process and is composed of five distinct loss components. These components consider the source image ($x_{src}$), the target image ($x_{trg}$), and the hair mask ($M$) as inputs. The combined loss function is defined as:
$\begin{gathered}\ell_{total}(x;x_{src},x_{trg},M)=\lambda_{1}\ell_{style}+\lambda_{2}\ell_{content}\\ +\lambda_{3}\ell_{mask}+\lambda_{4}\ell_{sem}+\lambda_{5}\ell_{rng},\end{gathered}$ (13)
@@
For $\ell_{style}$ and $\ell_{content}$, we utilize the style and content loss functions from DiffuseIT. We employ the [CLS] token matching loss using DINO ViT [6] to reflect semantic information in $x_{trg}$ and use keys of multi-head self-attention layers to preserve the content of $x_{src}$.
Additionally, to ensure hair preservation while translating scalp styles, we construct mask preservation loss function as:
$\ell_{mask}=\mbox{\small{LPIPS}}(x_{src}\odot M,\hat{x}_{0}(x_{t})\odot M)+||(x_{src}-\hat{x}_{0}(x_{t}))\odot M||_{2},$ (14)
where LPIPS denotes the learned perceptual image patch similarity metric [55].
We also include two other loss functions: $\ell_{rng}$, representing the squared spherical distance as proposed in [10], and $\ell_{sem}$, indicating the semantic divergence loss as outlined in [27].
Using this composite loss function, $\ell_{total}$, we guide the generation of the next sample step, $x_{t-1}$. To maintain hair information, we apply masking to the images:
$x_{t-1}=x_{t}\odot(1-M)+\Big{(}\hat{x}_{0}(x_{t})-\nabla_{x_{t}}\ell_{total}(\hat{x}_{0}(x_{t}))\Big{)}\odot M,$ (15)
where $\hat{x}_{0}(x_{t})$ is the estimation of the cleaned image derived from the sample $x_{t}$. This estimation is computed as follows:
@@
With this guide, we can translate the image style of the scalp without the need for additional training.
Classification strategy. Utilizing our DiffuseIT-M, as described above, we enhance our training set by translating randomly chosen images into ones with higher severity levels. We used random selection for the source images and weighted sampling for target images, where the likelihood of choosing an image was inversely proportional to the number of samples in its severity class.
For our classification task, we employed a pretrained backbone model and fine-tuned it with four Multi-Layer Perceptron (MLP) heads. Each head is associated with a specific loss function. The first head is designed to identify the presence of each scalp disease (i.e., dandruff, excess sebum, and erythema) without considering their severity. The remaining three heads are tasked with classifying the severity levels of each disease, categorized into four stages: good, mild, moderate, and severe. Consequently, our overall classification objective, denoted as $\ell_{cls}$, is expressed by the following formula:
$\ell_{cls}=\ell_{dis}+\ell_{dand}+\ell_{seb}+\ell_{ery},$ (17)
where $\ell_{dis}$ is the binary cross-entropy loss for the disease presence classification head, and $\ell_{dand}$, $\ell_{seb}$, and $\ell_{ery}$ represent the cross-entropy losses associated with determining the severity levels of each disease.

### v3

@@ Abstract
[…]
Keywords:
Scalp Disease Diagnosis Generative Data Augmentation.
@@ 1 Introduction
Scalp disorders are a widespread concern, with nearly $90\%$ of adults in the U.S. experiencing some form of condition [8]. Left unchecked, even seemingly mild scalp ailments can escalate into more serious outcomes, such as alopecia, underscoring the importance of timely intervention. Consequently, early diagnosis is crucial for preventing the progression of scalp-related diseases [23, 24], highlighting the need for advanced diagnostic approaches that are both efficient and accessible. Recognizing the importance of early detection, numerous studies have explored scalp disease diagnosis using microscopic scalp imagery [5, 14, 27].
Nevertheless, effectively diagnosing scalp disorders relies heavily on measuring critical features such as hair count and thickness, which demand precise hair segmentation. However, generating pixel-level hair annotations is costly and time-consuming, and no publicly available dataset provides such segmentation labels. The only major resource, AI-Hub [1], offers classification labels for scalp conditions but lacks segmentation annotations (see Section 3.1). Moreover, like many scalp image datasets, it suffers from data imbalance, especially for severe conditions, making it challenging to develop robust models.
To overcome these limitations, we propose ScalpVision, a comprehensive system for the in-depth assessment of scalp health. First, we achieve label-free hair segmentation by combining a naive segmentation model – trained on synthetic image-label pairs – with an automatic prompting module for the Segment Anything Model (SAM) [15], systematically generating positive and negative point prompts to enable accurate hair masks without manual labeling. Building on these masks, we then introduce DiffuseIT-M, a diffusion-based image-to-image translation framework that preserves hair details while altering scalp conditions. By generating diverse training samples, our method effectively mitigates data imbalance, ultimately leading to enhanced diagnostic performance for scalp diseases.
@@
Figure 1: ScalpVision pipeline overview: II is the original image, model $S$ generates the hair segmentation mask $\hat{M}$ using a pseudo-training set, $M_{\text{AP}}$ is the SAM-produced mask, and $M$ is the combined hair segmentation mask. The “Automatic Prompt” for refining segmentation comes from $\hat{M}$. $x_{src}$ and $x_{trg}$ are the source and target images, with $M$ as the mask image of $x_{src}$. The weighted image sum is denoted by $\odot$ and $D$ stands for DINO-ViT [4].
2 Method
@@
As illustrated in Figure 1, central to ScalpVision is a hair segmentation module (Section 2.1) and an image translation module for generating diverse scalp images to augment training datasets for scalp condition classification (Section 2.2).
@@ 2.1 Label-Free Hair Segmentation
For the precise diagnosis of scalp conditions, our initial step involves segmenting hair within microscopic scalp images. However, since most scalp condition datasets lack segmentation labels, supervised learning methods are not feasible.
Heuristic-driven pseudo-labeling. To address the absence of hair segmentation, we first generate pseudo labels for training our segmentation model (S as shown in Figure 1) using prior knowledge. With the intuition that the hair on the microscopic scalp images follows either a linear function or a power function, we generate synthetic images to effectively guide the model to learn hair patterns on the scalp images. For each disease condition, we randomly select one image representing each distinct severity level, extract three smaller patches from regions of the scalp with no visible hair, and draw curves to simulate hair patterns. Additionally, to simulate dandruff noise, circular white shapes are added to these patches but are not indicated in the pseudo masks, thus training the model to interpret them as noise. Examples of the pseudo images and masks are provided in the Appendix 0.C. We generate 3,0003{,}000 pseudo-images and corresponding pseudo mask labels, using them to train the U2-Net [25] which generates the binary mask, $\hat{M}=\hat{M}(i,j)\in\{0,1\}^{H\times W}$, where $H$ and $W$ are the height and width of the image, and $i\in[1,H]$, $j\in[1,W]$ denote pixel coordinates.
Input: Mask $\hat{M}$, bounding box size $n$, cross-shaped structuring element kernel
Output: Representative hair points from mask $\hat{\mathcal{C}}$
1
[…]
3
4while $\mathcal{H}_{\text{copy}}\neq 0$ do
5  Eroded $\leftarrow\texttt{MORPHOLOGY}\_\texttt{ERODE}(\mathcal{H}_{\text{copy}},\textit{kernel})$
6  $\textit{Dilated}\leftarrow\texttt{MORPHOLOGY}\_\texttt{DILATE}(\textit{Eroded},\textit{kernel})$
7  $\hat{K}\leftarrow\mathcal{H}_{\text{copy}}-\textit{Dilated}\,$; $\hat{\mathcal{H}}_{\text{skel}}\leftarrow\hat{\mathcal{H}}_{\text{skel}}\lor\hat{K}\,$; $\mathcal{H}_{\text{copy}}\leftarrow\textit{Eroded}$
8
9
10foreach $(x,y)\in\mathcal{H}_{\text{skel}}$ do
[…]
12
13
[…]
15foreach $(x_{1},y_{1},x_{2},y_{2})\in\hat{B}$ do
[…]
17
18
19return $\hat{\mathcal{C}}$
@@ 6  $\textit{Dilated}\leftarrow\texttt{MORPHOLOGY}\_\texttt{DILATE}(\textit{Eroded},\textit{kernel})$
Automatic prompting for SAM. To refine the hair segmentation mask $\hat{M}$, we utilize the foundation segmentation model, SAM [15], employing a point-prompting method to differentiate hair from scalp without additional training. However, selecting random points from $\hat{M}$ for positive prompts often led to suboptimal masks, mainly due to points near the edges of $\hat{M}$ confusing the SAM. Furthermore, the intrinsic randomness occasionally caused sampled points to coalesce within a confined region, thereby leading the SAM to segment only a limited subset of hairs. To address these issues, we developed an automatic prompting method, shown in Algorithm 1, that uniformly samples across $\hat{M}$ and uses the coarse segmentation mask $\hat{M}$ to guide the SAM with high confidence.
To extract the distinct features of the hair, we compute the skeletonized mask, $\hat{\mathcal{H}}_{\text{skel}}\in\{0,1\}^{H\times W}$, using morphological erosion and dilation following [34]. Then, we generate bounding boxes around each pixel in $\hat{\mathcal{H}}_{\text{skel}}$ with size $n\times n$ where we set $n=10$. These boxes undergo non-maximum suppression (NMS) to filter out the bounding boxes, denoted as $\hat{B}=\{\hat{b}_{j}\}_{j=1}^{k}$, where each box is defined by coordinates $(x_{\text{min}},y_{\text{min}},x_{\text{max}},y_{\text{max}})$. Following this, the mean points of the hair pixels, $\hat{\mathcal{C}}=\{\hat{c}_{j}\}_{j=1}^{k}$, in each bounding box $\hat{B}$ can be determined. For each $\hat{b}_{j}=(x_{1},y_{1},x_{2},y_{2})$, the mean point $\hat{c}_{j}=(\bar{x},\bar{y})$ is given by:
@@
Subsequently, we select positive point prompts for SAM from the calculated mean points $\hat{\mathcal{C}}$. For the negative point prompts, we utilize the inverse of the initial mask, specifically $1-\hat{M}$. These prompts, automatically generated, guide SAM in generating the binary segmentation mask, $M_{\text{AP}}\in\{0,1\}^{H\times W}$.
Mask ensemble. $M_{\text{AP}}$ and $\hat{M}$ complement each other with strengths and weaknesses. $\hat{M}$ is robust against noise like dandruff as it was trained using simulated noise. Meanwhile, $M_{\text{AP}}$, benefiting from SAM’s superior edge detection, excels in constructing a clear boundary between hair and scalp. Therefore, to make a robust hair mask, the final binary mask, $M$, is derived from $\hat{M}$ and $M_{\text{AP}}$ with the logical AND operation ($M=\hat{M}\land M_{\text{AP}}$), followed by a noisy region removal post-processing step with connected-component analysis.
@@ 2.2 Scalp Condition Classification
Accurately classifying scalp disease severity from microscopic images is difficult due to the rarity of extreme cases. To address this, we introduce DiffuseIT-M, a diffusion-based image translation model with mask guidance that transforms a source image into various scalp conditions while preserving hair content. Building on DiffuseIT [16] and incorporating an image editing technique inspired by blended diffusion [2], DiffuseIT-M enables robust augmentation of underrepresented classes for improved classification.
Image translation with mask guidance. To facilitate the transfer of scalp disease characteristics while preserving hair features in our model, we utilize a comprehensive loss function, $\ell_{total}$, that guides the reverse process and is composed of five distinct loss components. These components consider the source image ($x_{src}$), the target image ($x_{trg}$), and the hair mask ($M$) as inputs. The combined loss function is defined as:
$\begin{gathered}\ell_{total}\left(\scriptstyle{x;x_{src},x_{trg},M}\right)=\lambda_{1}\ell_{style}+\lambda_{2}\ell_{content}+\lambda_{3}\ell_{mask}+\lambda_{4}\ell_{sem}+\lambda_{5}\ell_{rng},\end{gathered}$ (2)
@@
For $\ell_{style}$ and $\ell_{content}$, we utilize the style and content loss functions from DiffuseIT, with further details provided in the Appendix 0.B. We employ the [CLS] token matching loss using DINO-ViT [4] to reflect semantic information in $x_{trg}$ and use keys of multi-head self-attention layers to preserve the content of $x_{src}$. Additionally, to ensure hair preservation while translating scalp styles, we construct a mask preservation loss function as:
[…]
where LPIPS denotes the learned perceptual image patch similarity metric [33] and $\hat{x}_{0}(x_{t})$ is the estimation of the cleaned image derived from the sample $x_{t}$:
@@
We also include two additional losses: $\ell_{rng}$, representing the squared spherical distance as proposed in [6], and $\ell_{sem}$, indicating the semantic divergence loss as outlined in [16]. Using this composite loss function, $\ell_{total}$, we guide the generation of the next sample step, $x_{t-1}$. To preserve hair details, we apply a masking approach:
$x_{t-1}\leftarrow x_{t}\odot M+\bigl{[}\hat{x}_{0}(x_{t})-\nabla{x_{t}}\ell_{total}\bigl{(}\hat{x}_{0}(x_{t})\bigr{)}\bigr{]}\odot(1-M).$ (5)
This method allows scalp style translation without extra training.
Classification strategy. Using DiffuseIT-M, we augment our training set by translating randomly chosen images into higher severity levels via weighted sampling, where selection probability is inversely proportional to the class’s size. The augmentation strategy and the corresponding results are provided in the Appendix 0.D. We fine-tune a pretrained backbone with four MLP heads, each tied to a specific loss. One head detects the presence of scalp diseases (dandruff, excess sebum, erythema), while the other three classify their severities (good, mild, moderate, severe). Our objective, $\ell_{cls}$, is the sum of four losses: $\ell_{dis}$ (binary cross-entropy for disease presence), and $\ell_{dand}$, $\ell_{seb}$, $\ell_{ery}$(cross-entropy for severity classification). This design enables simultaneous disease detection and severity assessment.
Figure 2: Data distribution of different severity within each scalp condition.
3 Experiments
3.1 Dataset
Despite the lack of publicly available datasets, we accessed a specialized dataset from AI-Hub [1] for classifying the severity of scalp dermatologic conditions. $\dagger$$\dagger\daggerThis$ dataset is provided by ‘The Open AI Dataset Project (AI-Hub, S. Korea)’ and is exempt from IRB approval as it does not contain any information that can identify individuals. The dataset is publicly accessible at https://aihub.or.kr. The dataset comprises 95,91095{,}910 images with a resolution of $640\times 480$ pixels from 20,00020{,}000 patients. It is split into 72,34272{,}342 training and 23,56823{,}568 test images, with 21,70321{,}703 from the training set used for validation. Dermatologists labeled each image for dandruff, excess sebum, and erythema, categorizing severity as good, mild, moderate, or severe. The dataset is heavily skewed toward good and mild cases (Fig. 2) and lacks segmentation labels. Thus, we manually annotated hair regions in $150$ test images to evaluate our hair segmentation methods.